# Finite presentability of twisted Brin–Thompson groups

### v1

@@ Introduction
One crucial aspect of our proof of Theorem A is a generalization of the so called Stein complex $X$ of $SV_{G}$ from [BZ22], to a family of subcomplexes $X(k)$ ($k\in\mathbb{N})$ that are significantly smaller. These may no longer be contractible like $X$, but are still simply connected for large enough $k$, and thus still useful for triyng to deduce finite presentability of $SV_{G}$.
@@
The Boone–Higman conjecture predicts that a finitely generated group has solvable word problem if and only if it embeds into a finitely presented simple group. The backward direction is easy, but the forward direction has been open for decades. This conjecture was posed by Boone and Higman in the early 1970’s [BH74], and has attracted renewed attention in recent years. Prominent examples of groups known to satisfy the conjecture include $\operatorname{GL}_{n}(\mathbb{Z})$ [Sco84], hence virtually special and virtually nilpotent groups, and also hyperbolic groups [BBMZa], and finitely presented or contracting self-similar groups [BBMZa, Zara, BM]. See [BBMZb] for more on the history and progress around this conjecture.
From now on, whenever we say that a group, “satisfies the Boone–Higman conjecture,” we are implicity saying that it has solvable word problem, and explicitly saying that it embeds into a finitely presented simple group. We should mention that we are not aware of any examples of groups for which the Boone–Higman conjecture is known to hold but the group does not embed into a group admitting an action of type (A). Thus, it is an interesting question whether embedding into a group admitting an action of type (A) could possibly be equivalent to embedding into a finitely presented simple group. In Section 5 we give some new examples of groups admitting actions of type (A), and discuss some general results related to the Boone–Higman conjecture.
@@ Acknowledgments
Many thanks are due to Jim Belk and Francesco Fournier-Facio; to Jim in particular for the idea of the quasi-retraction $\rho_{\kappa_{0}}$ in Section 2, and to Francesco in particular for the (idea for and) proof of Proposition 5.6. The author is supported by grant #635763 from the Simons Foundation.
@@ 1. Twisted Brin–Thompson groups and groupoids
More elegantly, this means $\tau_{\gamma}(\kappa)(\gamma s)=\kappa(s)$. The left/right, inverse/not inverse conventions here ensure that $\tau_{\gamma\gamma^{\prime}}=\tau_{\gamma}\tau_{\gamma^{\prime}}$. We will also say twist homeomorphism for anything of the form $h_{\psi}\circ\tau_{\gamma}\circ h_{\varphi}^{-1}$, so this is the twist homeomorphism from $B(\varphi)$ to $B(\psi)$ using $\gamma$.
@@ 2. Quasi-retractions
To see why these $\rho_{\kappa}$ produce a quasi-retraction from $SV_{G}$ to $\mathbb{Z}\wr_{S}G$, we need a series of lemmas.
@@
Choose the range partition for $g$ to equal the domain partition for $h$, say in the domain of $g$ we have $B(\varphi_{j})$, in the range of $g$ and domain of $h$ we have $B(\psi_{j})$, and in the range of $h$ we have $B(\chi_{j})$. Say $B(\varphi_{i})$ contains $\kappa$, and let $\gamma_{i}\in G$ be such that $g$ sends $B(\varphi_{i})$ to $B(\psi_{i})$ via the twist homeomorphism using $\gamma_{i}$. Similarly choose $\delta_{i}\in G$ such that $h$ takes $B(\psi_{i})$ to $B(\chi_{i})$ via the twist homeomorphism using $\delta_{i}$. Note that $hg$ takes $B(\varphi_{i})$ to $B(\chi_{i})$ via the twist homeomorphism using $\delta_{i}\gamma_{i}$. Now we compute
$\rho_{\kappa}(hg)$ $=\bigg{(}\big{(}r_{hg\kappa}^{s}(hg)-d_{\kappa}^{(\delta_{i}\gamma_{i})^{-1}s}(hg)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg\kappa}^{s}(h)-d_{\kappa}^{\gamma_{i}^{-1}\delta_{i}^{-1}s}(g)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg\kappa}^{s}(h)-d_{g\kappa}^{\delta_{i}^{-1}s}(h)+r_{g\kappa}^{\delta_{i}^{-1}s}(g)-d_{\kappa}^{\gamma_{i}^{-1}\delta_{i}^{-1}s}(g)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg\kappa}^{s}(h)-d_{g\kappa}^{\delta_{i}^{-1}s}(h)\big{)}_{s\in S},1\bigg{)}\bigg{(}\big{(}r_{g\kappa}^{\delta_{i}^{-1}s}(g)-d_{\kappa}^{\gamma_{i}^{-1}\delta_{i}^{-1}s}(g)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg\kappa}^{s}(h)-d_{g\kappa}^{\delta_{i}^{-1}s}(h)\big{)}_{s\in S},\delta_{i}\bigg{)}\bigg{(}\big{(}r_{g\kappa}^{s}(g)-d_{\kappa}^{\gamma_{i}^{-1}s}(g)\big{)}_{s\in S},\gamma_{i}\bigg{)}$
$=\rho_{g\kappa}(h)\rho_{\kappa}(g)\text{.}$
Note that we are still assuming $SV_{G}$ is finitely presented, hence finitely generated, so by [BZ22, Theorem A] we know $G$ is finitely generated and $S$ has finitely many $G$-orbits, which also tells us that $\mathbb{Z}\wr_{S}G$ is finitely generated.
@@
There exists a finite symmetric generating set $A$ for $SV_{G}$ such that the set $\{\rho_{\kappa}(a)\mid\kappa\in\mathfrak{C}^{S},a\in A\}$ is finite. In particular, there exists a finite symmetric generating set $B$ for $\mathbb{Z}\wr_{S}G$ such that $B$ contains $\{\rho_{\kappa}(a)\mid\kappa\in\mathfrak{C}^{S},a\in A\}$.
Here a subset of a group is symmetric if it is closed under taking inverses.
@@
The second statement is immediate from the first, so we just need to prove the first. Let us build up to a generating set for $SV_{G}$. First, let $\iota_{\varnothing}\colon G\to SV_{G}$ be the monomorphism sending $\gamma$ to the twist homeomorphism $\tau_{\gamma}$. Next, for $s\in S$, let $\iota_{1}^{s}\colon G\to SV_{G}$ be the monomorphism sending $\gamma$ to the homeomorphism $x_{s}^{-1}(\operatorname{id}\oplus\tau_{\gamma})x_{s}$, so intuitively this applies the twist homeomorphism $\tau_{\gamma}$ to the “second half” of $\mathfrak{C}^{S}$ in the dimension $s$. Finally, for any $T\subseteq S$ consider the embedding $TV\hookrightarrow SV$ by writing $\mathfrak{C}^{S}=\mathfrak{C}^{T}\times\mathfrak{C}^{S\setminus T}$ and having $TV$ act trivially on the $\mathfrak{C}^{S\setminus T}$ factor. Write $TV^{(S)}$ for this subgroup of $SV$.
By the proof of [BZ22, Theorem A] (in Section 3 of that paper), $SV_{G}$ is generated by $\iota_{\varnothing}(G)$, $\iota_{1}^{s_{0}}(G)$ for any choice of $s_{0}\in S$, and finitely many $\{s,t\}V^{(S)}$ for $s,t\in S$. Note that for any $\gamma\in G$ and any $\kappa$, $\rho_{\kappa}$ sends $\iota_{\varnothing}(\gamma)$ to $((0)_{s\in S},\gamma)$ and sends $\iota_{1}^{s_{0}}(G)$ to either $((0)_{s\in S},1_{G})$ or $((0)_{s\in S},\gamma)$. Thus, we can choose any finite symmetric generating sets we like for $\iota_{\varnothing}(G)$ and $\iota_{1}^{s_{0}}(G)$, and their contributions to the set of all $\rho_{\kappa}(a)$ will be finite.
It remains to choose a finite symmetric generating set for each $\{s,t\}V^{(S)}$ that contributes only finitely many $\rho_{\kappa}(a)$. In other words, we are now free to assume without loss of generality that $S=\{1,2\}$ or $S=\{1\}$ (for the $s\neq t$ and $s=t$ cases respectively) and $G=\{1\}$, so in particular we can view $\rho_{\kappa}$ as a function from $2V$ to $\mathbb{Z}^{2}$, or from $V$ to $\mathbb{Z}$ depending on whether $s=t$. A generating set for $2V$ is given in [Bri04, Subsection 5.3]. This generating set also appears in [BC10, Section 2], along with convenient pictures of the generators in [BC10, Figure 4]. It is clear, especially from the pictures, that for any such generator or inverse generator $a$, and any $\kappa\in\mathfrak{C}^{2}$, we have $\rho_{\kappa}(a)\in\{(0,0),(1,0),(0,1),(1,1)\}$. The $s=t$ case works the same, using the obvious analogous generating set for $V$. ∎
@@
First we need to construct a function $\zeta\colon\mathbb{Z}\wr_{S}G\to SV_{G}$. Embed $\mathbb{Z}$ into $V$ by sending $1$ to some element that acts by sending the dyadic brick $B(0)$ in $\mathfrak{C}$ to $B(00)$ via $0\kappa\mapsto 00\kappa$. Extend this to $\bigoplus_{S}\mathbb{Z}\to\bigoplus_{S}V$. Now observe that $\bigoplus_{S}V$ embeds into $SV$ by having the copy of $V$ corresponding to $s\in S$ act on the factor of $\mathfrak{C}$ in $\mathfrak{C}^{S}$ corresponding to $s$, and fixing all other coordinates. Finally, $G$ embeds into $SV_{G}$ via the map $\iota_{\varnothing}$, i.e., as a group of global coordinate permutations. Clearly this copy of $G$ normalizes the copy of $\bigoplus_{S}V$, and acts by permuting coordinates, so together they form a copy of $V\wr_{S}G$. Putting everything together we get a monomorphism $\zeta\colon\mathbb{Z}\wr_{S}G\to SV_{G}$.
Now we claim that $\rho_{\kappa_{0}}\circ\zeta$ is the identity on $\mathbb{Z}\wr_{S}G$. Note that every element in the image of $\zeta$ fixes $\kappa_{0}$, so restricted to this image, $\rho_{\kappa_{0}}$ is a homomorphism by Lemma 2.5. Thus, it suffices to check that $\rho_{\kappa_{0}}\circ\zeta$ is the identity on both the $\bigoplus_{S}\mathbb{Z}$ and $G$ factors individually, or even just on some generating sets thereof. On the $G$ factor, $\zeta(\gamma)=\iota_{\varnothing}(\gamma)$, which $\rho_{\kappa_{0}}$ sends to $\gamma$ as desired. Now let $(z_{s})_{s\in S}\in\bigoplus_{S}\mathbb{Z}$, and assume that it comes from the standard generating set, so $z_{s}=1$ for some $s\in S$ and $z_{t}=0$ for all $t\neq s$. Let $\varphi\colon S\to\{0,1\}^{*}$ send $s$ to 00 and all other $t$ to $\varnothing$. Restricted to $B(\varphi)$, the element $h=\zeta((z_{s})_{s\in S},1_{G})$ acts in the $s$ dimension by adding a zero and acts trivially in all other dimensions. Since $B(\varphi)$ contains $\kappa_{0}$, and it is possible to find for $h$ a domain partition into dyadic bricks including $B(\varphi)$ and a range partition into dyadic bricks including the image of $B(\varphi)$ under $h$, this shows that $r_{\kappa_{0}}^{s}(h)-d_{\kappa_{0}}^{s}(h)=1$ and $r_{\kappa_{0}}^{t}(h)-d_{\kappa_{0}}^{t}(h)=0$ for all $t\neq s$.We conclude that $\rho_{\kappa_{0}}(h)=((z_{s})_{s\in S},1_{G})$ as desired.
@@
This also shows that if $SV_{G}$ is of type $\operatorname{F}_{n}$ then so is $\mathbb{Z}\wr_{S}G$, which imposes restrictions on $G$ and $S$, thanks to [BdCK15, Theorem B]. Namely, we must have that $G$ is of type $\operatorname{F}_{n}$, each $\operatorname{Stab}_{G}(T)$ for $T\subseteq S$ with $|T|<n$ is of type $\operatorname{F}_{n-|T|}$, and there are finitely many orbits of $n$-element subsets of $S$. This is conjectured to be an if and only if in [BZ22, Conjecture H], so it is worth remarking here that our quasi-retract argument in this section has now proven the “only if” direction of this conjecture (and the main result of the present paper is that the “if” direction is also true for $n=2$).
@@ 3. Finite presentability from actions with bad edge stabilizers
Let us first recall how a group presentation falls out of a cellular action on a simply connected complex. In [Bro84] the complex can be any CW-complex and the action is allowed to stabilize cells without fixing them. For simplicity, here we will follow [Arm88] and assume the complex is simplicial and the action is rigid, meaning the stabilizer of each cell equals its fixer. See also [Put11] for a nice simplification in the case when the orbit space is $2$-connected (which is unfortunately not the case for the situations we will care about later).
@@
Let $K$ be a simply connected simplicial complex. View the $1$-skeleton $K^{(1)}$ as a (simplicial) graph. We will assume this graph is directed, so each edge $e$ has an origin $o(e)$ and terminus $t(e)$, and we write $\overline{e}$ for $e$ with the opposite orientation, so $o(\overline{e})=t(e)$ and $t(\overline{e})=o(e)$. Let $\Gamma$ be a group with an orientation preserving action on $K$, i.e., for any edge $e$ and any $\gamma\in\Gamma$ we have $o(\gamma.e)=\gamma.o(e)$ and $t(\gamma.e)=\gamma.t(e)$. Let $M$ be a maximal tree in the orbit space $\Gamma\backslash K^{(1)}$, and let $T$ be a choice of lift of $M$ to a subtree of $K^{(1)}$. Note that the vertices of $T$ form a set of representatives of the $\Gamma$-orbits of vertices of $K$. For each edge $f$ of $(\Gamma\backslash K^{(1)})\setminus M$, choose a lift $\widetilde{f}$ of $f$ in $K^{(1)}$ such that one endpoint of $\widetilde{f}$ lies in $T$, call it the origin $o(\widetilde{f})$. The other endpoint of $\widetilde{f}$, its terminus $t(\widetilde{f})$, necessarily lies outside $T$. Let $z_{f}$ be the vertex of $T$ that shares a $\Gamma$-orbit with $t(\widetilde{f})$, and choose some $\gamma_{f}\in\Gamma$ such that $\gamma_{f}.z_{f}=t(\widetilde{f})$. Note that we are free to choose $\gamma_{f}$ to be any element we like, as long as it satisfies $\gamma_{f}.z_{f}=t(\widetilde{f})$. We impose the rule that we always choose $\gamma_{\overline{f}}=\gamma_{f}^{-1}$. For each edge $f$ of $M$, set $\gamma_{f}=1$.

### v2

@@ Introduction
As a remark, the sufficient conditions in [BZ22] to get $SV_{G}$ to be finitely presented, assuming $G$ is, required that $\operatorname{Stab}_{G}(T)$ be finitely presented for all finite $T\subseteq S$, and that for each $k\in\mathbb{N}$ there are finitely many $G$-orbits of $k$-element subsets. In [Zarb] finite presentability of the stabilizers was relaxed to finite generation. The “if” part of our main theorem here is therefore a significant improvement on these existing results, since now one only needs to understand stabilizers of single points, and orbits of pairs.
We should also point out one crucial aspect of our proof of Theorem A, which is a generalization of the so called Stein complex $X$ of $SV_{G}$ from [BZ22], to a family of subcomplexes $X(k)$ ($k\in\mathbb{N})$ that are significantly smaller. These may no longer be contractible like $X$, but are still simply connected for large enough $k$, and thus still useful for trying to deduce finite presentability of $SV_{G}$.
@@
The Boone–Higman conjecture predicts that a finitely generated group has solvable word problem if and only if it embeds into a finitely presented simple group. The backward direction is easy, but the forward direction has been open for decades. This conjecture was posed by Boone and Higman in the early 1970’s [BH74], and has attracted renewed attention in recent years. Prominent examples of groups known to satisfy the conjecture include $\operatorname{GL}_{n}(\mathbb{Z})$ [Sco84], hence virtually special and virtually nilpotent groups, and also hyperbolic groups [BBMZa], finitely presented or contracting self-similar groups [BBMZa, Zara, BM], and Baumslag–Solitar groups and (finite rank free)-by-cyclic groups [BLIW]. See [BBMZb] for more on the history and progress around this conjecture.
From now on, whenever we say that a group, “satisfies the Boone–Higman conjecture,” we are implicity saying that it has solvable word problem, and explicitly saying that it embeds into a finitely presented simple group. We should mention that we are not aware of any examples of groups for which the Boone–Higman conjecture is known to hold but the group does not embed into a group admitting an action of type (A). Thus, it is an interesting question (see Question 5.13) whether embedding into a group admitting an action of type (A) could possibly be equivalent to embedding into a finitely presented simple group. In Section 5 we give some new examples of groups admitting actions of type (A), and discuss some general results related to the Boone–Higman conjecture.
@@ Acknowledgments
Many thanks are due to Jim Belk and Francesco Fournier-Facio; to Jim in particular for the idea of the quasi-retraction $\rho_{\kappa_{0}}$ in Section 2, and to Francesco in particular for the (idea for and) proof of Proposition 5.6. I am also grateful to the referee for many helpful comments and suggestions. The author is supported by grant #635763 from the Simons Foundation.
@@ 1. Twisted Brin–Thompson groups and groupoids
More elegantly, this means $\tau_{\gamma}(\kappa)(\gamma s)=\kappa(s)$. As in [BZ22], we view the action of $G$ on $S$ as a left action, and the left/right, inverse/not inverse conventions here ensure that $\tau_{\gamma\gamma^{\prime}}=\tau_{\gamma}\tau_{\gamma^{\prime}}$. We will also say twist homeomorphism for anything of the form $h_{\psi}\circ\tau_{\gamma}\circ h_{\varphi}^{-1}$, so this is the twist homeomorphism from $B(\varphi)$ to $B(\psi)$ using $\gamma$.
@@ 2. Quasi-retractions
To see why these $\rho_{\kappa}$ are useful for getting a quasi-retraction from $SV_{G}$ to $\mathbb{Z}\wr_{S}G$, we need a series of lemmas.
@@
Choose the range partition for $g$ to equal the domain partition for $h$; say in the domain of $g$ we have $B(\varphi_{j})$, in the range of $g$ and domain of $h$ we have $B(\psi_{j})$, and in the range of $h$ we have $B(\chi_{j})$. Say $B(\varphi_{i})$ contains $\kappa$, and let $\gamma_{i}\in G$ be such that $g$ sends $B(\varphi_{i})$ to $B(\psi_{i})$ via the twist homeomorphism using $\gamma_{i}$. Similarly choose $\delta_{i}\in G$ such that $h$ takes $B(\psi_{i})$ to $B(\chi_{i})$ via the twist homeomorphism using $\delta_{i}$. Note that $hg$ takes $B(\varphi_{i})$ to $B(\chi_{i})$ via the twist homeomorphism using $\delta_{i}\gamma_{i}$. Now we compute
$\rho_{\kappa}(hg)$ $=\bigg{(}\big{(}r_{hg(\kappa)}^{s}(hg)-d_{\kappa}^{(\delta_{i}\gamma_{i})^{-1}s}(hg)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg(\kappa)}^{s}(h)-d_{\kappa}^{\gamma_{i}^{-1}\delta_{i}^{-1}s}(g)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg(\kappa)}^{s}(h)-d_{g(\kappa)}^{\delta_{i}^{-1}s}(h)+r_{g(\kappa)}^{\delta_{i}^{-1}s}(g)-d_{\kappa}^{\gamma_{i}^{-1}\delta_{i}^{-1}s}(g)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg(\kappa)}^{s}(h)-d_{g(\kappa)}^{\delta_{i}^{-1}s}(h)\big{)}_{s\in S},1\bigg{)}\bigg{(}\big{(}r_{g(\kappa)}^{\delta_{i}^{-1}s}(g)-d_{\kappa}^{\gamma_{i}^{-1}\delta_{i}^{-1}s}(g)\big{)}_{s\in S},\delta_{i}\gamma_{i}\bigg{)}$
$=\bigg{(}\big{(}r_{hg(\kappa)}^{s}(h)-d_{g(\kappa)}^{\delta_{i}^{-1}s}(h)\big{)}_{s\in S},\delta_{i}\bigg{)}\bigg{(}\big{(}r_{g(\kappa)}^{s}(g)-d_{\kappa}^{\gamma_{i}^{-1}s}(g)\big{)}_{s\in S},\gamma_{i}\bigg{)}$
$=\rho_{g(\kappa)}(h)\rho_{\kappa}(g)\text{.}$
Note that we are still assuming $SV_{G}$ is finitely presented, hence finitely generated, so by [BZ22, Theorem A] we know $G$ is finitely generated and $S$ has finitely many $G$-orbits, which also tells us that $\mathbb{Z}\wr_{S}G$ is finitely generated (see for example [dC06, Proposition 2.1]). At this point we fix some finite symmetric generating set $A$ for $SV_{G}$. (Here a subset of a group is symmetric if it is closed under taking inverses.)
@@
There exists a finite symmetric generating set $B$ for $\mathbb{Z}\wr_{S}G$ such that $B$ contains $\{\rho_{\kappa}(a)\mid\kappa\in\mathfrak{C}^{S},a\in A\}$.
@@
Since we already know $\mathbb{Z}\wr_{S}G$ is finitely generated, all we have to do is prove that $\{\rho_{\kappa}(a)\mid\kappa\in\mathfrak{C}^{S},a\in A\}$ is finite. Indeed, we claim that for any $h\in SV_{G}$ the set $\{\rho_{\kappa}(h)\mid\kappa\in\mathfrak{C}^{S}\}$ is finite. This is because, fixing a domain and range partition that encode $h$, if $\kappa$ and $\kappa^{\prime}$ share a domain block then by construction we have $\rho_{\kappa}(h)=\rho_{\kappa^{\prime}}(h)$. Since our fixed domain partition has finitely many blocks, $\{\rho_{\kappa}(h)\mid\kappa\in\mathfrak{C}^{S}\}$ is finite. Now since $A$ is finite, $\{\rho_{\kappa}(a)\mid\kappa\in\mathfrak{C}^{S},a\in A\}$ is finite. ∎
@@
First we need to construct a function $\zeta\colon\mathbb{Z}\wr_{S}G\to SV_{G}$. Embed $\mathbb{Z}$ into $V$ by sending $1$ to some element that acts by sending the dyadic brick $B(0)$ in $\mathfrak{C}$ to $B(00)$ via $0\kappa\mapsto 00\kappa$. Extend this to $\bigoplus_{S}\mathbb{Z}\to\bigoplus_{S}V$. Now observe that $\bigoplus_{S}V$ embeds into $SV$ by having the copy of $V$ corresponding to $s\in S$ act on the factor of $\mathfrak{C}$ in $\mathfrak{C}^{S}$ corresponding to $s$, and fixing all other coordinates. Finally, $G$ embeds into $SV_{G}$ as a group of global coordinate permutations; following [BZ22] we write $\iota_{\varnothing}\colon G\to SV_{G}$ for this embedding. Clearly this copy of $G$ normalizes the copy of $\bigoplus_{S}V$, and acts by permuting coordinates, so together they form a copy of $V\wr_{S}G$. Putting everything together we get a monomorphism $\zeta\colon\mathbb{Z}\wr_{S}G\to SV_{G}$.
Now we claim that $\rho_{\kappa_{0}}\circ\zeta$ is the identity on $\mathbb{Z}\wr_{S}G$. Note that every element in the image of $\zeta$ fixes $\kappa_{0}$, so restricted to this image, $\rho_{\kappa_{0}}$ is a homomorphism by Lemma 2.5. Thus, it suffices to check that $\rho_{\kappa_{0}}\circ\zeta$ is the identity on both the $\bigoplus_{S}\mathbb{Z}$ and $G$ factors individually, or even just on some generating sets thereof. On the $G$ factor, $\zeta(\gamma)=\iota_{\varnothing}(\gamma)$, which $\rho_{\kappa_{0}}$ sends to $\gamma$ as desired. Now let $(z_{s})_{s\in S}\in\bigoplus_{S}\mathbb{Z}$, and assume that it comes from the standard generating set, so $z_{s}=1$ for some $s\in S$ and $z_{t}=0$ for all $t\neq s$. Let $\varphi\colon S\to\{0,1\}^{*}$ send $s$ to 00 and all other $t$ to $\varnothing$. Restricted to $B(\varphi)$, the element $h=\zeta((z_{s})_{s\in S},1_{G})$ acts in the $s$ dimension by prepending a zero and acts trivially in all other dimensions. Since $B(\varphi)$ contains $\kappa_{0}$, and it is possible to find for $h$ a domain partition into dyadic bricks including $B(\varphi)$ and a range partition into dyadic bricks including the image of $B(\varphi)$ under $h$, this shows that $r_{\kappa_{0}}^{s}(h)-d_{\kappa_{0}}^{s}(h)=1$ and $r_{\kappa_{0}}^{t}(h)-d_{\kappa_{0}}^{t}(h)=0$ for all $t\neq s$.We conclude that $\rho_{\kappa_{0}}(h)=((z_{s})_{s\in S},1_{G})$ as desired.
@@
This also shows that if $SV_{G}$ is of type $\operatorname{F}_{n}$ then so is $\mathbb{Z}\wr_{S}G$, which imposes restrictions on $G$ and $S$, thanks to [BdCK15, Theorem B]. Namely, we must have that $G$ is of type $\operatorname{F}_{n}$, each $\operatorname{Stab}_{G}(T)$ for $T\subseteq S$ with $|T|<n$ is of type $\operatorname{F}_{n-|T|}$, and there are finitely many orbits of $n$-element subsets of $S$. Let us call the action of $G$ on $S$ in this case an action of type (An), so type (A2) means type (A). This is conjectured to be an if and only if in [BZ22, Conjecture H], so it is worth remarking here that our quasi-retract argument in this section has now proven the “only if” direction of this conjecture (and the main result of the present paper is that the “if” direction is also true for $n=2$). It remains open whether the action of $G$ on $S$ being of type (An) implies $SV_{G}$ is of type $\operatorname{F}_{n}$.
@@ 3. Finite presentability from actions with bad edge stabilizers
Let us first recall how a group presentation falls out of a cellular action on a simply connected complex. In [Bro84] the complex can be any CW-complex and the action is allowed to stabilize cells without fixing them. For simplicity, here we will follow [Arm88] and assume the complex is simplicial and the action is rigid, meaning the stabilizer of each cell equals its fixer. See also [Put11] for a nice simplification in the case when the orbit space is $2$-connected (which is unfortunately not the case for the situations we will care about later). All this coming discussion before Proposition 3.1 is taken directly from [Arm88].
@@
Let $K$ be a simply connected simplicial complex. View the $1$-skeleton $K^{(1)}$ as a (simplicial) graph. As in [Arm88], following Serre [Ser77], each edge $e$ can be made into a pair of directed edges by specifying which endpoint is the origin $o(e)$ and which is the terminus $t(e)$. Write $\overline{e}$ for $e$ with the opposite orientation, so $o(\overline{e})=t(e)$ and $t(\overline{e})=o(e)$, and both $e$ and $\overline{e}$ are viewed as (directed) edges in $K^{(1)}$. Let $\Gamma$ be a group with an orientation preserving action on $K$, i.e., for any edge $e$ and any $\gamma\in\Gamma$ we have $o(\gamma.e)=\gamma.o(e)$ and $t(\gamma.e)=\gamma.t(e)$. Let $M$ be a maximal tree in the orbit space $\Gamma\backslash K^{(1)}$, and let $T$ be a choice of lift of $M$ to a subtree of $K^{(1)}$. Note that the vertices of $T$ form a set of representatives of the $\Gamma$-orbits of vertices of $K$. For each edge $f$ of $(\Gamma\backslash K^{(1)})\setminus M$, choose a lift $\widetilde{f}$ of $f$ in $K^{(1)}$ such that one endpoint of $\widetilde{f}$ lies in $T$, call it the origin $o(\widetilde{f})$. The other endpoint of $\widetilde{f}$, its terminus $t(\widetilde{f})$, necessarily lies outside $T$. Let $z_{f}$ be the vertex of $T$ that shares a $\Gamma$-orbit with $t(\widetilde{f})$, and choose some $\gamma_{f}\in\Gamma$ such that $\gamma_{f}.z_{f}=t(\widetilde{f})$. Note that for each $f$ we are free to choose $\gamma_{f}$ to be any element we like, as long as it satisfies $\gamma_{f}.z_{f}=t(\widetilde{f})$. We will also always choose the lift $\widetilde{(\overline{f})}$ of $\overline{f}$ to have $z_{f}$ specifically as its origin, so consequently $z_{\overline{f}}$ is the origin of $\widetilde{f}$, and we can take $\gamma_{\overline{f}}=\gamma_{f}^{-1}$. For each edge $f$ of $M$, set $\gamma_{f}=1$.